A triangular region has a perimeter of 91 meters. The first side is two-thirds of the second side. The third side is 13 meters shorter than the second side. What are the lengths of the three sides of the triangular region?

Answer by checkley79(3341)

Let x be the second side.
Let 2x/3 for the first side.
Let x-13 be the third side.
91=x+2x/3+x-13
91=(3x+2x+3x)/3-13
91+13=8x/3
104=8x/3
8x=104*3
8x=312
x=312/8
x=39 length of the second side
2*39/3=78/3=26 length of the first side.
39-13=26 length of the third side.
Proof:
91=39+26+26
91=91